# SmallPlan: Leverage Small Language Models for Sequential Path Planning with LLM-Guided Distillation and Reinforcement Learning from Simulation Feedback

### v5

@@ II Related Works
Definition of SLMs. Categorizations of SLMs vary in their contexts [24]. For examples, [25] defines SLMs as models with fewer than 1 billion params, while [26, 27, 28] consider the term SLMs relative to the larger counterparts. For mobile devices, typically possessing around 6GB of memory, [29] suggests that SLMs contain fewer than 10 billion params. In this work, all employed SLMs are capped at 3 billion params and require less than 4GB memory for inference.
@@ III Problem Statement and Scenario Set-up
Resource Constraints. We assume $\mathcal{M}$ runs real-time on edge devices, with 6GB of memory. In our work, the SLMs are capped at 3 billion parameters with less than 4GB of memory for inference.
@@ IV-B2 Reinforcement Learning with Simulation Feedback
MoMaLLM [14] GPT-4o 97.14 65.92 18.01 0.39
Pretrained Qwen-2.5 61.33 46.59 12.00 3.05
Pretrained Llama-3 53.68 38.02 9.84 4.16
Pretrained Phi4-mini 70.86 41.57 17.84 3.39
Ours Qwen-2.5 83.43 48.07 19.87 1.04
Ours Llama-3 93.14 66.10 17.65 1.22
Ours Phi4-mini 96.57 70.41 16.13 0.49
IV-B3 SmallPlan Training Strategy
@@
In the first stage, the candidate SLM is supervised fine-tuned with a small set of response samples from the teacher LLM. Pretrained SLMs often fail to produce outputs in a consistent format, even with well-crafted instruction prompts, making it difficult to parse their responses into executable commands. This step thus helps the SLM learns the correct output format from the LLMs. This few-shot SFT strategy also fits with the LoRA setup as it helps the SLMs to highlight the important parameters for the target task. To speed up this stage, we prepare these sample subsets offline beforehand and train SFT with these samples.
In the second stage, we train the SLM using an interleaved combination of RL and SFT. At runtime, we collect environment data including detected objects, their distances, and their relations from the scene graph and Voronoi graph, then parse this information into the prompt template to build a complete prompt as input for LLM and SLM to response. Here, SFT acts as a distillation mechanism for transferring reasoning behavior from the LLM, while RL - using our $r_{\text{action}}$ reward - adjusts the model’s planning behavior and controls task-related factors. This stage can also be configured to use only SFT, as distillation alone can sometimes be enough to achieve strong performance. In our experiments in Section V, we compare different training strategies to evaluate their effectiveness. Our framework, however, strongly discourages removing the SFT and relying solely on RL, as this would prevent the SLM from effectively learning from the LLM.
@@ V-B4 Out-of-domain Analysis
Fig. 3 evaluates the logical reasoning capabilities of the SLMs on out-of-domain tasks beyond path planning. These tasks involve different types of sequential decision-making, allowing us to assess general reasoning ability. Our trained SLMs perform on par with, and sometimes exceed, their pretrained versions, as reflected by the consistent SR with only minor differences. This suggests that distillation from a superior LLM can help retain the general reasoning ability of smaller models, even across diverse domains. Given the complexity of cross-domain adaptation, we leave further investigation to future research works. This experiment merely highlights that our fine-tuned SLMs maintain strong generalization and are not overfitting to path planning tasks. The goal here is not to compare different SLMs against each other on their success rates, but to compare each trained SLM with its own pretrained version to demonstrate that our SLMs retain their general reasoning capability.
@@ V-B5 Hyper-parameter Selection and Prompting
We try various hyper-parameter configurations during training, including RL rewards and the number of training epochs. Based on our experience, selecting these settings is largely a trial-and-error process, as no standardized method can guarantee optimal results. By default, we use $r_{\text{success}}=5$, $\lambda_{\text{success}}=0.3$, $\lambda_{\text{explore}}=0.1$, $\lambda_{\text{efficiency}}=0.3$, and $\lambda_{\text{format}}=0.1$. For penalization of Dist. traveled, we increase the efficiency weight to $\lambda_{\text{efficiency}}=0.6$. The learning rate is set at $1\times 10^{-5}$ for all training. Meanwhile, the number of epochs depends heavily on the SLM architecture, as different models vary in learning capacity and convergence. We use 4 epochs for Phi-4-mini and Llama-3, and 6 epochs for Qwen-2.5. Simply increasing or decreasing the number of epochs does not consistently improve performance. We tested multiple settings and selected the best-performers. Our models achieved their best performance with 3 to 6 epochs. Further tuning efforts can help to find optimal hyper-parameters and further improve the SLMs’ performance.

### v6

@@ II Related Works
Definition of SLMs. Categorizations of SLMs varies across contexts [24]. For example, [25] defines SLMs as models with fewer than 1 billion parameters, while [26, 27, 28] consider the term SLMs relative to their larger counterparts. For mobile devices, typically possessing around 6GB of memory, [29] suggests SLMs should have fewer than 10 billion parameters. NVIDIA [30] similarly considers models under 10 billion parameters as SLMs. In this work, we study the SLMs with at most 4 billion parameters and require less than 4GB of memory for inference.
@@ III Problem Statement and Scenario Set-up
Resource Constraints. We assume $\mathcal{M}$ runs real-time on edge devices, with 6GB of memory. In our work, the SLMs are capped at 4 billion parameters with less than 4GB of memory for inference.
@@ IV-B2 Reinforcement Learning with Simulation Feedback
MoMaLLM [14] GPT-4o 97.14 65.92 18.01 0.39
Pretrained Qwen-2.5 61.33 46.59 12.00 3.05
Pretrained Llama-3 53.68 38.02 9.84 4.16
Pretrained Phi4-mini 70.86 41.57 17.84 3.39
Ours Qwen-2.5 83.43 48.07 19.87 1.04
Ours Llama-3 93.14 66.10 17.65 1.22
Ours Phi4-mini 96.57 70.41 16.13 0.49
IV-B3 SmallPlan Training Strategy
In the first stage, the candidate SLM is supervised fine-tuned with a small set of response samples from the teacher LLM. Pretrained SLMs often fail to produce outputs in a consistent format, even with well-crafted instruction prompts, making it difficult to parse their responses into executable commands. This step thus helps the SLM learns the correct output format from the LLMs. This few-shot SFT strategy also fits with the LoRA setup as it helps the SLMs to highlight the important parameters for the target task. To speed up this stage, we collect the LLM responses and store them offline, and train SFT with these samples.
In the second stage, we train the SLM using an interleaved combination of RL and SFT. At runtime, we collect environment data including detected objects, their distances, and their relations from the scene graph and Voronoi graph, then parse this information into the prompt template to build a complete prompt as input for LLM and SLM to response. Here, SFT acts as a distillation mechanism for transferring reasoning behavior from the LLM, while RL - using our $r_{\text{action}}$ reward - adjusts the model’s planning behavior and controls task-related factors. This stage can also be configured to use only SFT, as distillation alone can sometimes be enough to achieve strong performance. In our experiments in Section V, we compare different training strategies, RL-SFT and SFT only, to evaluate their effectiveness. Our framework, however, strongly discourages removing the SFT and relying solely on RL, as this would prevent the SLM from effectively learning from the LLM.
@@ V-B4 Out-of-domain Analysis
Fig. 3 evaluates the logical reasoning capabilities of the SLMs on out-of-domain tasks beyond path planning. These tasks involve different types of sequential decision-making, allowing us to assess general reasoning ability. Our trained SLMs perform on par with, and sometimes exceed, their pretrained versions, as reflected by the consistent SR with only minor differences. This suggests that distillation from a superior LLM can help retain the general reasoning ability of smaller models, even across diverse domains. Given the complexity of cross-domain adaptation, we leave further investigation to future research. This experiment merely highlights that our trained SLMs maintain strong generalization and are not overfitting to path planning tasks. The goal here is not to compare different SLMs against each other on their success rates, but to compare each trained SLM with its pretrained version to show that our SLMs retain their general reasoning capability.
@@ V-B5 Hyper-parameter Selection and Prompting
We try various hyper-parameter settings during training, including RL rewards and the number of training epochs. Based on our experience, selecting these settings is largely a trial-and-error process, as no standardized method can guarantee optimal results. We use $r_{\text{success}}=5$, $\lambda_{\text{action\_executable}}=0.3$, $\lambda_{\text{explore}}=0.1$, $\lambda_{\text{efficiency}}=0.3$, and $\lambda_{\text{format}}=0.1$. For penalization of Dist. traveled, we increase the efficiency weight to $\lambda_{\text{efficiency}}=0.6$. Learning rate is fixed at $1\times 10^{-5}$ for all training. Meanwhile, the number of epochs depends heavily on the SLM architectures, as they vary in learning capacity and convergence. For stage 1 (Algorithm 1), we fine-tune with 3 epochs using 500 offline data samples for all SLMs. For stage 2, we use 4 epochs for Phi-4-mini and Llama-3, and 6 epochs for Qwen-2.5. Simply increasing or decreasing the number of epochs does not consistently improve performance. We tested multiple settings and selected the best-performers. Our models perform best with 3 to 6 epochs. Further tuning efforts can help to find optimal hyper-parameters and further improve the SLMs’ performance.